# HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?

### v2

@@ 4.3 T-mAP Hyperparameters
T-mAP has two hyperparameters: the maximum allowed time delta $\delta$ and evaluation horizon $H$. We set $\delta$ twice the cost of the OTD because when the model predicts an incorrect label, OTD removes the prediction and adds the ground truth event with the total cost equal to $2C$. The horizon $H$ must be larger than $\delta$ to evaluate timestamp quality adequately. Therefore, depending on the dataset, we select $H$ to be approximately 3-7 times larger than $\delta$, ensuring the horizon captures an average of 6-15 events. The empirical study of T-mAP hyperparameters is presented in Figure 4. As shown, the chosen $\delta$ parameter is positioned to the right of the initial slope of the parameter-quality curve, ensuring an optimal balance between time granularity and task difficulty.
Figure 4: T-mAP dependency on the $\delta$ parameter. The dashed line indicates the selected value. Results for the StackOverflow and Amazon datasets are provided in the Appendix E.
5 HoTPP Benchmark
@@
The HoTPP benchmark integrates data preprocessing, training, and evaluation in a single toolbox. Unlike previous benchmarks, HoTPP introduces the novel T-mAP metric for long-horizon prediction. HoTPP differs from prior MTPP benchmarks by including simple rule-based baselines and next-k models, simultaneously predicting multiple future events. The HoTPP benchmark is designed with a focus on simplicity, extensibility, evaluation stability, reproducibility, and computational efficiency.
@@
Simplicity and Extensibility. The benchmark code is organized into a clear structure, separating the core library, dataset-specific scripts, and configuration files. New methods can be integrated at various levels, including the configuration file, model architecture, loss function, metric, and training module. Core components are reusable through the Hydra configuration library (Yadan, 2019). For example, HoTPP can apply the HYPRO rescoring method to backbone models such as IFTPP, RMTPP, NHP, and ODE. Another example is our Next-K implementation, which can be applied to IFTTP or RMTPP.
@@
Evaluation Stability. In the MTPP domain, many datasets contain only a few thousand sequences for training and evaluation. For example, the StackOverflow test set includes 401 sequences, Retweet contains 1956, and Amazon has 1851 sequences for testing. Previous long-horizon evaluation pipelines (Xue et al., 2022; 2023) made predictions only at the end of each sequence, resulting in a limited number of predictions and reduced evaluation stability. To address this limitation, we evaluate long-horizon predictions at multiple intermediate points.
@@
Reproducibility. The HoTPP benchmark ensures reproducibility at multiple levels. First, we use the PytorchLightning library (Falcon & The PyTorch Lightning team, 2019) for training with a specified random seed and report multi-seed evaluation results. Second, data preprocessing is carefully designed to ensure datasets are constructed reproducibly. Finally, we specify the environment in a Dockerfile.
Computational Efficiency. Some methods are particularly slow, especially during autoregressive inference on large datasets. Straightforward evaluation can take several hours on a single Nvidia V100 GPU. To accelerate computation, we optimized the training and inference pipelines in two ways. First, we implemented an efficient RNN that reuses computations during parallel autoregressive inference from multiple starting positions. Second, we developed highly optimized versions of the thinning algorithm used for sampling in NHP and continuous-time neural architectures (NHP and ODE), achieving up to a 4x performance improvement compared to the official implementations. These optimizations allowed us to conduct the first large-scale evaluations of algorithms such as NHP, AttNHP, ODE, and HYPRO on datasets like Transactions and MIMIC-IV. Additionally, we provide the first open-source CUDA implementation of the batched linear assignment solver, significantly enhancing the applicability of the proposed T-mAP metric.
A detailed description of the benchmark can be found in Appendix C, with further details on HoTPP’s computational performance in Appendix C.5.
@@
Table 1: Evaluation results. The best result is shown in bold. The mean and standard deviation of each metric computed during five runs with different random seeds are reported.

### v3

@@ 3.2.3. T-mAP Hyperparameters
T-mAP has two hyperparameters: the maximum allowed time delta $\delta$ and evaluation horizon $H$. We set $\delta$ twice the cost of the OTD because when the model predicts an incorrect label, OTD removes the prediction and adds the ground truth event with the total cost equal to $2C$. The horizon $H$ must be larger than $\delta$ to evaluate timestamp quality adequately. Therefore, depending on the dataset, we select $H$ to be approximately 3-7 times larger than $\delta$, ensuring the horizon captures an average of 6-15 events. The empirical study of T-mAP hyperparameters is presented in Fig. 4. As shown, the chosen $\delta$ parameter is positioned to the right of the initial slope of the parameter-quality curve, ensuring an optimal balance between time granularity and task difficulty.
4. HoTPP Benchmark
@@
The HoTPP benchmark integrates data preprocessing, training, and evaluation in a single toolbox. Unlike previous benchmarks, HoTPP introduces the novel T-mAP metric for long-horizon prediction. HoTPP also differs from prior MTPP benchmarks by including simple rule-based baselines and long-horizon prediction methods that simultaneously predict multiple future events, such as HYPRO (Xue et al., 2022) and Diffusion (Zhou et al., 2025). The HoTPP benchmark is designed to focus on simplicity, extensibility, evaluation stability, reproducibility, and computational efficiency.
@@
Simplicity and Extensibility. The benchmark code is organized into a clear structure, separating the core library, dataset-specific scripts, and configuration files. New methods can be integrated at various levels, including the configuration file, model architecture, loss function, metric, and training module. Core components are reusable through the Hydra configuration library (Yadan, 2019). For example, HoTPP can apply the HYPRO rescoring method to backbone models such as IFTPP, RMTPP, NHP, or ODE. Furthermore, each method’s backbone architecture can be easily switched between RNNs and Transformers. Methods such as IFTPP, RMTPP, and Diffusion are compatible with most HuggingFace models (Wolf et al., 2020).
@@
Evaluation Stability. In the MTPP domain, many datasets contain only a few thousand sequences for training and evaluation. For example, the StackOverflow test set includes 401 sequences, Retweet contains 1956, and Amazon has 1851 sequences for testing. This possesses a unique challenge in the long-horizon forecasting task. Previous long-horizon evaluation pipelines (Xue et al., 2022, 2023) made predictions only at the end of each sequence, resulting in a limited number of measurements and a low evaluation stability. To address this limitation, we evaluate long-horizon predictions at multiple intermediate points, decreasing the variance of evaluation metrics.
Table 1. Datasets statistics
Dataset # Sequences # Events Mean Length Mean Horizon Mean Duration # Classes OTD T-mAP Length Steps Cost $\delta$ Horizon Transactions 50k 43.7M 875 9.0 719 203 5 1 2 7 MIMIC-IV 120k 2.4M 19.7 6.6 503 34 5 2 4 28 Retweet 23k 1.3M 56.4 14.7 1805 3 10 15 30 180 Amazon 9k 403K 43.6 14.8 22.1 16 5 1 2 10 StackOverflow 2k 138K 64.2 12.0 55.3 22 10 1 2 10
@@
Reproducibility. The HoTPP benchmark ensures reproducibility at multiple levels. First, we use the PytorchLightning library (Falcon and The PyTorch Lightning team, 2019) for training with a specified random seed and report multi-seed evaluation results. Second, data preprocessing is carefully designed to ensure datasets are constructed reproducibly. Finally, we specify the environment in a Dockerfile.
@@
Table 2. Evaluation results. The best result is shown in bold. The mean and standard deviation of each metric computed during five runs with different random seeds are reported.